# TRAM: Global Trajectory and Motion of 3D Humans from in-the-wild Videos

### v1

@@ Abstract
We propose TRAM, a two-stage method to reconstruct a human’s global trajectory and motion from in-the-wild videos. TRAM robustifies SLAM to recover the camera motion in the presence of dynamic humans and uses the scene background to derive the motion scale. Using the recovered camera as a metric-scale reference frame, we introduce a video transformer model (VIMO) to regress the kinematic body motion of a human. By composing the two motions, we achieve accurate recovery of 3D humans in the world space, reducing global motion errors by 60% from prior work. https://yufu-wang.github.io/tram4d/
@@ 1 Introduction
By robustifying and metrifying the camera estimation, and reconstructing accurate human body motion in the camera frame, our two-stage approach recovers the global trajectory and motion (TRAM) of 3D humans from videos, with a 60% error reduction in global motion estimation from the best published result.
Contributions. (i) We propose a general method TRAM, that recovers human trajectory and motion from in-the-wild videos with a large improvement over prior works. (ii) We demonstrate that the human trajectory can be inferred from SLAM, and provide technical solutions to make monocular SLAM robust and metric-scale in the presence of dynamic humans. (iii) We propose the video transformer model VIMO that builds on top of a large pre-trained image-based model, and demonstrate that this scalable design achieves state-of-the-art reconstruction performance.
@@ 3.4 Video Transformer for Human Motion
Human pose and shape regression networks can now reconstruct 3D people from a single image in the general case, but videos provide temporal constraints that can further disambiguate the problem and help reconstruct motions that look natural and smooth. In this work, we propose VIdeo Transformer for Human MOtion (VIMO) to accomplish this goal.
A recent success recipe in vision is to finetune a large pre-trained model for downstream tasks. HMR2.0 [17] demonstrates state-of-the-art human pose and shape reconstruction by building on top of a pre-trained ViT-H [20, 91] and scaling up its training data and compute. Our design philosophy for VIMO is therefore twofold: we aim to utilize the rich knowledge in a large pre-trained model, and make flexible and scalable architecture design whose performance can continue to improve with compute and data.
Toward this goal, we turn the image-based HMR2.0 into a video model, by adding two temporal transformers. We finetune them with the original transformer decoder on video data. As shown in Fig. 3, the two temporal modules use the same encoder-only transformer architecture but distinctively propagate temporal information in the image domain and the human motion domain.
@@ 4 Experiments
For human trajectory evaluation, we slice a sequence into 100-frame segments and evaluate 3D joint error after aligning the first two frames (W-MPJPE100) or the entire segment (WA-MPJPE100[93]. In addition, we follow WHAM [70] and evaluate the error of the entire trajectory after aligning the first frame, with root translation error (RTE), root orientation error (ROE), and egocentric root velocity error (ERVE).
For camera trajectory evaluation, we follow SLAM literature and evaluate absolute trajectory error (ATE) [75], which performs Procrustes with scaling to align the estimation with ground truth before computing error. To evaluate the accuracy of our scale estimation, we evaluate ATE using our estimated scale (ATE-S) [35].
@@ 4.1 Comparison on Camera Trajectory Recovery
Models    PA–MPJPE W-MPJPE100 WA-MPJPE100 RTE ROE ERVE
GLAMR [94]  56.0 756.1 286.2 16.7 74.9 18.0
TRACE [77]  58.0 2244.9 544.1 18.9 72.7 370.7
SLAHMR [93]  61.5 807.4 336.9 13.8 67.9 19.7
WHAM (w/ DROID ) [70]  41.9 439.2 166.1 8.4 36.0 14.7
TRAM  38.1 222.4 76.4 3.1 13.6 10.3
Table 3: Evaluation of human global trajectory and motions. RTE is in $m$, ROE is in $deg$, ERVE is in $mm/\text{frame}$, and the other pose metrics are in $mm$.
@@ 4.3 Comparison on Human Body Motion Reconstruction
Figure 6: Human motion and trajectory: WHAM   vs Ours  . We produce more accurate motion and tracks that generalize to diverse terrain and motion complexity. Results from WHAM are generated with the ground truth gyro.
@@
HMR2.0 [17]  44.4 69.8 82.2 18.1    60.6 98.0 120.3 19.8
BEDLAM-CLIFF [6]  43.0 66.9 78.5 31.0    60.6 98.0 111.6 36.1
temporal TCMR [11]  52.7 86.5 101.4 6.0    79.6 127.6 147.9 5.3
VIBE [33]  51.9 82.9 98.4 18.5    81.4 125.9 146.8 26.6
MPS-Net [87]  52.1 84.3 99.0 6.5    81.3 123.1 138.4 6.2
GLoT [69]  50.6 80.7 96.4 6.0    78.8 119.7 138.4 5.4
GLAMR [94]  51.1 8.0    73.5 113.6 133.4 32.9
TRACE [77]  50.9 79.1 95.4 28.6    70.9 109.9 127.4 25.5
WHAM-B [70]  37.2 59.4 71.0 6.9    48.8 80.7 93.7 5.9
@@ 5 Conclusions
We presented TRAM, a two-stage method to recover the global human trajectory and body motion from in-the-wild videos with moving cameras. TRAM is efficient and accurate, improving global human reconstruction by large margins. We also introduced VIMO, a video transformer model for regressing the local human body motion. VIMO is simple and scalable, and outperforms prior models in different benchmarks.
While we make progress in this problem, much can be improved. Depth prediction is not guaranteed to be accurate in all cases. We see evidence that the rendered depth of the SMPL model can potentially help correct the bias in depth prediction. We include this experiment in the supplementary. An improved depth model [92, 67] will also improve the accuracy. Another interesting direction is to reason about spatial relationships with large vision-language models [9]. We have shown that disentangling the camera and human reconstruction improves robustness, but it can be beneficial to perform global optimization using our results as initialization. It’s an interesting direction to study how to jointly update human, camera, and scene reconstruction [35].
@@
Supplementary Material
F Additional Experiments
F.1 Naive Combination: DROID + ZoeDepth
Figure 7: Camera trajectory. Top row: examples where the naive combination achieves comparable results. Bottom row: naive combination leads to large error.
[…]
F.2 Using SMPL Depth to Improve ZoeDepth
Figure 8: Improving depth prediction. We render depth from SMPL reconstruction, and use this depth map to correct any shift or scale biases in the metric depth prediction. We scale and shift the depth prediction so that its human region would align with the rendered SMPL depth.
  EMDB 2 (ATE-S)
Scale Estimation using    Short(5) Medium(10) Long(10) Average
ZoeDepth    0.48 0.62 0.78 0.66
ZoeDepth + shift correction    0.37 0.36 1.41 0.78
ZoeDepth + scale-shift correction    0.35 0.37 1.43 0.79
Table 6: Camera scale estimation. Using SMPL depth rendering to correct scale and shift in depth prediction produces mixed results.
We have shown that ZoeDepth prediction is not always accurate. Particularly, there could be shift and scale biases. In such cases, the metric depth prediction can be regarded as affine-invariant depth prediction. If there are objects of known depth in the image, we can use them to correct the shift and scale. Can we use human mesh reconstruction to help correct the biases? Specifically, could we estimate shift and scale variables $s$ and $t$ to correct depth prediction $\hat{D}=s*D+t$?
Figure 8 illustrates this approach. We solve for the scale and shift correction by aligning the human region in the depth prediction to the rendered depth from SMPL reconstruction, through energy minimization similar to the robust optimization in Sec 3.3 of the main text. We report the quantitative results in Table 6. We observe mixed results: it improves scale estimation in some sequences but decreases accuracy in others. Specifically, we observe that it decreases ATE-S (better) by $20\%$ in 10/25 sequences but increases ATE-S (worse) by $20\%$ in 5/25 sequences. The average ATE-S is slightly worse, because worse cases happen to be long sequences, so a small error in scale estimation could lead to a much higher translation error.
The effectiveness of this approach is also influenced by the accuracy of the mesh reconstruction. If the predicted human shape is more accurate, it will be more effective. Inaccurate shape prediction (e.g., the predicted human being is taller than the ground truth) will produce inaccurate depth rendering.
G Implementation Details
Figure 9: Architecture of VIMO. Left: the detailed architecture of VIMO, with the yellow blocks denoting the new temporal components. Right: the architecture of the two temporal transformers.
G.1 Architecture
We show a more detailed view of the VIMO architecture in Figure 9. VIMO interleaves spatial and temporal modules. Both temporal transformers have 6 layers and 4 multi-head attention. The first temporal transformer (image domain) has an embedding dimension of 512, while the second temporal transformer (motion domain) has an embedding dimension of 384.
G.2 Datasets
We use 3DPW, Human3.6M, and BEDLAM to train our video transformer VIMO. We evaluate on 3DPW and EMDB. 3DPW is an in-the-wild dataset providing ground truth 3D pose annotations acquired with IMU and videos. 2D and 3D joints are generated from the pose annotation. Human3.6M is an indoor multi-view dataset with 2D and 3D joint annotation. Additionally, we use SMPL recovered using MoSH for this dataset. BEDLAM is a large synthetic dataset rendered with Unreal Engine 5 and SMPL. Therefore, it has the most accurate SMPL pose and shape. EMDB is an in-the-wild dataset with accurate SMPL and trajectory annotations recovered with electromagnetic sensors.
During training, we sample sequences of 16 frames from the three datasets. There are about 1.3k sequences from 3DPW, 19k from Human3.6M, and 305k from BEDLAM (30fps). So we sample sequences unequally from each dataset to guarantee a good mix of real and synthetic data, with the following ratio: [3DPW: 16.5%, Human3.6M: 16.5%, BEDLAM: 67%].
G.3 ORB-SLAM2
We use the open source ORB-SLAM2 implementation released by the authors in https://github.com/raulmur/ORB_SLAM2. For the masked evaluation, we first process the images in the dataset by setting all pixels within the human masks to a value of 255. We run the entire ORB-SLAM2 pipeline including camera tracking, point reconstruction, and loop closure. We specify our configuration based on the default monocular SLAM parameters for the TUM RGB-D dataset provided in the code and increase the number of features detected at each frame to 4,000. Additionally, because the EMDB videos sometimes demonstrate low contrast with a uniform background, we slightly lowered the minimum fast feature threshold per image patch (more details in the ORB-SLAM2 configuration documentation). Despite these efforts, we still observed tracking failures due to a fast-moving camera as well as textureless background regions. Compared to ORB-SLAM2, DROID performed better in handling these areas because they do not rely on distinctive sparse features; instead, DROID uses optical flow to guide correspondence, which shares the benefits of low-texture texture handling with direct SLAM methods. For loop closure, we use the bag-of-words vocabulary provided in the official repository.
G.4 Training
Acceleration. The training of video models is costly. Because the backbone is often frozen, previous methods pre-compute the features output by the backbone and use them as input to finetune the upper layers (including new components). While this approach reduces forward time, it is impossible to apply data augmentation. To address this issue, we do not pre-compute features but use two other methods: pre-cropping images and half-precision backbone inference. The datasets provide high-resolution images which could take longer to load, crop, and resize. Pre-cropping the images and saving them as crops reduces loading time. Using crops has a different disadvantage: data augmentation such as random rotation and scaling will produce black borders. While it could potentially reduce the effectiveness (it’s not clear the extent), it is still better than no augmentation. Secondly, we use half-precision inference for the backbone, which reduces forward time. Since we do not finetune the backbone, using half-precision will not affect the training.
Data Augmentation. We apply standard data augmentation including rotation, scaling, horizontal flipping, color jittering, and occlusion. For video model training, all augmentations except occlusion are applied consistently in the same sequence. For example, the same degree of random rotation should be applied for all 16 frames of a sequence. However, each frame has an independent and equal chance of having occlusion augmentation.

### v2

@@ Abstract
We propose TRAM, a two-stage method to reconstruct a human’s global trajectory and motion from in-the-wild videos. TRAM robustifies SLAM to recover the camera motion in the presence of dynamic humans and uses the scene background to derive the motion scale. Using the recovered camera as a metric-scale reference frame, we introduce a video transformer model (VIMO) to regress the kinematic body motion of a human. By composing the two motions, we achieve accurate recovery of 3D humans in the world space, reducing global motion errors by a large margin from prior work. https://yufu-wang.github.io/tram4d/
@@ 1 Introduction
Our method takes a scene-centric point of view in estimating human movements, versus recent methods with a human-centric point of view using motion priors. By solely using the scene background to estimate metric scale camera motion, and always reconstructing humans in the camera frame, our two-stage approach recovers the global trajectory and motion (TRAM) of 3D humans from videos, with a large error reduction in global trajectory from the best published results.
Contributions. (i) We propose a general method TRAM, that recovers human trajectory and motion from in-the-wild videos with a large improvement over prior works. (ii) We demonstrate that the human trajectory can be inferred from SLAM, and provide technical solutions to make monocular SLAM robust and metric-scale in the presence of dynamic humans. (iii) We propose the video transformer model VIMO that builds on top of a large pre-trained image-based model, and demonstrate that this scalable design achieves state-of-the-art pose and shape reconstruction.
@@ 3.4 Video Transformer for Human Motion
Human pose and shape regression networks can now reconstruct 3D people from a single image in the general case, but videos provide temporal constraints to help reconstruct motions that look natural and smooth. In this work, we propose VIdeo Transformer for Human MOtion (VIMO) to accomplish this goal.
A recent success recipe in vision is to finetune a large pre-trained model for downstream tasks. HMR2.0 [17] demonstrates state-of-the-art reconstruction by building on top of a pre-trained ViT-H [20, 93] and scaling up its training data and compute. Our design philosophy for VIMO is therefore twofold: we aim to utilize the rich knowledge in a large pre-trained model, and make flexible and scalable architecture design whose performance can scale with data.
Toward this goal, we turn the image-based HMR2.0 into a video model by adding two temporal transformers, as shown in Fig. 3. HMR2.0 uses a ViT-H to extract image features and a standard transformer decoder with a zero token to cross-attend the image features and regress the final outputs. The two new temporal transformers from VIMO use the encoder-only architecture but distinctively propagate temporal information in the image domain and the human motion domain.
@@ 4 Experiments
[…]
For camera trajectory evaluation, we follow SLAM literature and evaluate absolute trajectory error (ATE) [76], which performs rigid alignment with scaling to align the camera trajectory with ground truth before computing error. To evaluate the accuracy of our scale estimation, we also evaluate ATE using our estimated scale (ATE-S) [36].
@@ 4.1 Comparison on Camera Trajectory Recovery
Models    PA–MPJPE WA-MPJPE100 W-MPJPE100 RTE ERVE
TRACE [78]  58.0 529.0 1702.3 17.7 370.7
GLAMR [96]  56.0 280.8 726.6 11.4 18.0
SLAHMR [95]  61.5 326.9 776.1 10.2 19.7
WHAM (w/ DROID ) [70]  38.2 133.3 343.9 4.6 14.7
TRAM  38.1 76.4 222.4 1.4 10.3
[…]
@@ 4.3 Comparison on Human Body Motion Reconstruction
Figure 6: Human motion and trajectory: WHAM   vs Ours  . We produce more accurate motion and tracks that generalize to diverse terrain and motion complexity. WHAM’s results are from the version that uses ground truth gyro.
@@
HMR2.0 [17]  44.4 69.8 82.2 18.1    60.7 98.3 120.8 19.9
ReFit [88]  40.5 65.3 75.1 18.5    58.6 88.0 104.5 20.7
temporal TCMR [11]  52.7 86.5 101.4 6.0    79.8 127.7 150.2 5.3
VIBE [34]  51.9 82.9 98.4 18.5    81.6 126.1 149.9 26.5
MPS-Net [89]  52.1 84.3 99.0 6.5    81.4 123.3 143.9 6.2
GLoT [69]  50.6 80.7 96.4 6.0    79.1 119.9 140.8 5.4
GLAMR [96]  51.1 8.0    73.8 113.8 134.9 33.0
TRACE [78]  50.9 79.1 95.4 28.6    71.5 110.0 129.6 25.5
WHAM (ViT) [70]  35.9 57.8 68.7 6.6    50.4 79.7 94.4 5.3
@@
4.4 Limitations
While TRAM works well on datasets, we detail the following challenges for future research. Firstly, SLAM’s dependence on known focal length limits its applicability in many in-the-wild cases. Future research will need to address the estimation of focal length during bundle adjustment [8, 72]. Secondly, depth estimation tends to be less accurate with extreme focal length, and methods that consider focal length will be beneficial [23]. Lastly, our method follows a strict separation of camera and human motion but a joint optimization in the end could be beneficial [36]. Joint optimization with proper physics priors will improve implausible movements such as foot sliding and penetration [81].
@@ 5 Conclusions
We presented TRAM, a new two-stage method to recover the global human trajectory and body motion from in-the-wild videos with moving cameras. TRAM is efficient and accurate, improving global human reconstruction by large margins. We also introduced VIMO, a video transformer model for regressing the local human body motion. VIMO is simple and scalable, and outperforms prior models in different pose and shape reconstruction benchmarks.